A graduated cylinder contains 155mL of water. a 15.0 g piece of iron and a 20.0 g piece of lead are added. What is the new water level?


1
Expert's answer
2019-09-02T03:39:42-0400

density of iron: d(Fe)=7.9 g/cm3

density of lead: d(Pb)=11.3 g/cm3


"V(add)=V(Fe)+V(Pb)=m(Fe)\/d(Fe)+m(Pb)\/d(Pb)=15.0\/7.9+20.0\/11.3=3.7mL"


"V(new)=V(water)+V(add)=155+3.7=158.7mL"